Suppose the weather forecast calls for a "60% chance of rain" each day for the next 3 days. What is the probability that it will NOT rain during the next 3 days? Assume that whether it rains each day is an independent event. Round answer to 2 places after the decimal point.
Answer #1

Required probability = P(it will not rain during the next 3 days)

= P(it does not rain on Day 1) x P(it does not rain on Day 2) x P(it does not rain on Day 3)

= (1-0.6)(1-0.6)(1-0.6)

= 0.064 0.06